# Parallel Batch Scheduling With Incompatible Job Families Via Constraint Programming

### v1

@@ II Literature Review
In another recent development in the AM industry, Cakici et al. [10] presented a synchronized CP model that replaces binary variables with interval variables, resulting in improved performance over the AS formulation. However, they consider only a single family of compatible jobs. Furthermore, their model leverages optional interval variables in combination with a global synchronization constraint to ensure that jobs within the same batch start and finish together. While this approach is computationally efficient, it can lead to synchronization inconsistencies due to the optional nature of the variables involved in the synchronization process. Specifically, their model does not explicitly enforce the presence of the required batches when synchronizing jobs, which may result in unintended outcomes. This paper addresses this limitation by introducing a constraint that guarantees the presence of a batch whenever a job is assigned to it, thereby ensuring consistent synchronization across all jobs and batches. Moreover, this paper extends their model to handle multiple incompatible job families and impose additional constraints to accommodate batch size requirements.
@@ III Contributions
As evidenced in the previous section, most of the existing CP models for p-batch only deal with a single compatible family [8, 9, 10]. Among these, the only model that synchronizes the jobs does not enforce the presence of a batch when jobs are assigned to it [10], resulting in an inconsistent schedule. Moreover, the existing CP model that considers multiple incompatible families fails to synchronize the jobs within their batch, allowing interruptions mid-process [18].
This paper addresses these gaps in the literature by proposing three new CP models that consider multiple incompatible job families and directly tackle the interruptions in their formulation while synchronizing the jobs within their batch: (i) An Assign-and-Schedule (AS) model inspired by Nascimento et al. [8, 9] that handles the assignment of jobs to batches with binary variables, and the sequencing of batches on machines using CP scheduling logic; (ii) A Synchronized (S) model inspired by Cakici et al. [10] that replaces the binary assignment with CP scheduling logic, and guarantees the correct job synchronization by ensuring the “presence of batches” that have assigned jobs; (iii) and a new Redundant Synchronized (RS) model that includes redundant variables and constraints that help the search process and improves computational and quality performance. Furthermore, the paper explores the impact of symmetry-breaking constraints in the search process. The computational experiments demonstrate the superiority of the RS model over existing models.
@@ V The Existing CP Model
Model 1 presents the CP model proposed by Ham et al. [18] to schedule incompatible jobs in the diffusion area. Equation (1) defines an interval variable $x_{j}$ of the form $[a,a+p_{j})$ (i.e., it has a duration of exactly $p_{j}$ units of time), which represents job $j\in\mathcal{J}$. Equation (1) defines an optional interval variable $x_{jm}$, also of size $p_{j}$, that represents the option of processing job $j$ on machine $m\in\mathcal{M}$. The presence of the interval $x_{jm}$ is optional since it is allowed to take the value $\perp$, which indicates its absence in the solution. These optional interval variables allow the model to decide which machine should process the jobs. Equation (1) defines a function ${\mathfrak{f}}_{m}$, which is the unique state of the machine $m$ at every point in time. Equation (1) defines a cumulative function ${\mathscr{l}}_{mf}$, which represents the load on machine $m$ of jobs of the family $f\in\mathcal{F}$. It uses the pulse function, which receives two inputs: An interval $x_{jm}$ and an integer value $s_{j}$. This pulse function tells the cumulative function to increment its value in $s_{j}$ units during the interval $x_{jm}$, if present. Hence, the function ${\mathscr{l}}_{mf}$ accumulates the sizes of overlapping jobs of family $f\in\mathcal{F}$ on machine $m$ at every point in time.
@@
$x_{j}$$\in\{[a,a+p_{j}):a\in\mathbb{Z}\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J};$ $(1)$
@@ VI-B The Synchronized Model
$x_{jm}^{b}$$\in\{[a,a+p_{j})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M},b\in\mathcal{B}_{f_{j}};$ $(22)$
$y_{bm}$$\in\{[a,a+p^{f^{b}})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(23)$
@@
Model 3 presents the Synchronized (S) model. It is based on the CP model by Cakici et al. [10] for AM processes, which deals with a single compatible family; this paper extends their formulation to handle multiple incompatible families and include the necessary constraints to ensure a consistent synchronization of the jobs. Equation (3) defines an optional interval variable $x_{jm}^{b}$, of size $p_{j}$, that represents the option of job $j$ being batched in batch $b\in\mathcal{B}_{f_{j}}$ and processed on machine $m\in\mathcal{M}$. Equation (3) defines an optional interval variable $y_{bm}$ of size $p^{f^{b}}$ that represents the option of batch $b\in\mathcal{B}$ being processed on machine $m$. Equation (3) defines an interval sequence variable $\varphi_{m}$ that represents the sequence of batches scheduled on machine $m$. Equation (3) defines a cumulative function ${\mathscr{l}}_{m}^{b}$ that represents the load on machine $m$ due to the jobs in batch $b$.
@@
Objective function (3) minimizes the TWCT. Constraints (3) assign each job to exactly one batch on exactly one machine. Constraints (3) allow each batch to be processed in at most one machine. Constraints (3) use the synchronize global constraint, which takes two inputs: an interval variable $v=x_{jm}^{b}$ and a set of interval variables $B=\set{y_{bm}}$. If $v$ is present, this constraint attempts to synchronize it with any intervals in $B$ that are also present [38], i.e., $\texttt{presenceOf}(v)=1\implies\texttt{startOf}(v)=\texttt{startOf}(b)\wedge\texttt{endOf}(v)=\texttt{endOf}(b),\forall\leavevmode\nobreak\ b\in B$ s.t. $\texttt{presenceOf}(b)=1$. In practice, when the interval variable representing the job being assigned to a batch on a machine (i.e., $x_{jm}^{b}$) is part of the solution, the synchronize constraint attempts to align its start and end times with those of the batch on the machine (i.e., $y_{bm}$), if it is also present. However, the synchronize constraint only ensures synchronization if both interval variables are present; it does not inherently enforce the presence of $y_{bm}$ when $x_{jm}^{b}$ is scheduled. This subtle issue, where the presence of $x_{jm}^{b}$ should logically imply the presence of $y_{bm}$, is not addressed in the original model by Cakici et al. [10], leading to synchronization inconsistencies. To resolve this, constraints (3) introduce an explicit implication constraint, which ensures that, whenever $x_{jm}^{b}$ is included in the solution, the corresponding batch interval $y_{bm}$ is also present, thereby guaranteeing consistent synchronization as intended. Constraints (3) use the noOverlap global constraint to ensure that the present interval variables in the machine sequence do not overlap, guaranteeing at most one batch being processed at a time on each machine. Finally, constraints (3) guarantee that the load of the machines when processing each job in its batch is within the size limits of the batch’s family.
@@ VI-C The Redundant Synchronized Model
$x_{j}$$\in\{[a,a+p_{j})\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J};$ $(33)$
$x_{jm}$$\in\{[a,a+p_{j})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(34)$
$x_{jm}^{b}$$\in\{[a,a+p_{j})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M},b\in\mathcal{B}_{f_{j}};$ $(35)$
$y_{b}$$\in\{[a,a+p^{f^{b}})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B};$ $(36)$
$y_{bm}$$\in\{[a,a+p^{f^{b}})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(37)$
@@
$V_{jm}^{b}$ =$=\{y_{b},y_{bm},x_{jm}\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M},b\in\mathcal{B}_{f_{j}};$ $(39)$
@@ VIII Concluding remarks
This paper addressed the incompatible case of parallel batch scheduling, where compatible jobs belong to the same family, and jobs of different families cannot be processed together in the same batch. In the context of the diffusion area of the SM process, batches cannot be interrupted mid-process and jobs must be synchronized within their batch. The NS model by Ham et al. [18] for this problem does not synchronize the jobs within their batch, allowing interruptions mid-process. To remedy this limitation, the paper presented three new CP models, including two adaptions of existing models and a novel RS model. The AS model of Nascimento et al. [8] was extended to handle multiple incompatible job families and satisfy the batch size requirements. The S model of Cakici et al. [10] was also extended to the incompatible case while ensuring a consistent synchronization of jobs within their batches. Finally, the paper proposed a new RS model that exclusively uses global constraints and adds redundancy to the problem structure. The computational experiments demonstrate that the proposed RS model dominates all the other models: it finds solutions to all the instances, proves optimality in the largest amount of instances, produces solutions that are significantly better than other models, especially in the largest instances and achieves the best gaps in the least amount of time. Current research includes using CP for serial batching, and future research includes extending the efficient arc-flow formulations for compatible p-batch [16, 29, 28, 30] to the incompatible case and embedding the proposed formulations in larger scheduling contexts [40, 41, 42].

### v2

@@ II Literature Review
In another recent development in the AM industry, Cakici et al. [10] presented a synchronized CP model that replaces binary variables with interval variables, resulting in improved performance over the AS formulation. However, they consider only a single family of compatible jobs. This paper extends their model to handle multiple incompatible job families and impose additional constraints to accommodate batch size requirements. Furthermore, their model leverages optional interval variables in combination with a global synchronization constraint to ensure that jobs within the same batch start and finish together. A challenge faced in the AM industry is that the duration of the batches depends on the physical size of the parts being 3D printed in them. This results in only knowing the batch duration after the jobs are assigned to them. Their model defines a constraint that ensures the correct duration of the batches while simultaneously guaranteeing their presence when needed. In contrast, in the diffusion context studied in this paper, batches can be predefined to only group jobs of a specific job family, allowing to know a priori their duration. Hence, such constraint that defines the batch duration is not necessary, but the presence of the correct batches is still required. This paper uses a different approach to do so and includes an explicit implication constraint instead.
@@ III Contributions
As evidenced in the previous section, most of the existing CP models for p-batch only deal with a single compatible family [8, 9, 10]. Among these, the only model that synchronizes the jobs relies on a constraint specifically designed to ensure the correct batch duration in the AM context, which does not directly translate to the diffusion problem [10]. Moreover, the existing CP model that considers multiple incompatible families fails to synchronize the jobs within their batch, allowing interruptions mid-process [18].
This paper addresses these gaps in the literature by proposing three new CP models that consider multiple incompatible job families and directly tackle the interruptions in their formulation while synchronizing the jobs within their batch: (i) An Assign-and-Schedule (AS) model inspired by Nascimento et al. [8, 9] that handles the assignment of jobs to batches with binary variables, and the sequencing of batches on machines using CP scheduling logic; (ii) A Synchronized (S) model inspired by Cakici et al. [10] that replaces the binary assignment with CP scheduling logic, and guarantees the presence of the required batches with an implication constraint; (iii) and a new Redundant Synchronized (RS) model that includes redundant variables and constraints that help the search process and improves computational and quality performance. Furthermore, the paper explores the impact of symmetry-breaking constraints in the search process. The computational experiments demonstrate the superiority of the RS model over existing models.
@@ V The Existing CP Model
Model 1 presents the CP model proposed by Ham et al. [18] to schedule incompatible jobs in the diffusion area. Equation (1) defines an interval variable $x_{j}$ of the form $[a,a+p_{j})$ (i.e., it has a duration of exactly $p_{j}$ units of time), which represents job $j\in\mathcal{J}$. Note that this interval can only start after the job’s release time. Equation (1) defines an optional interval variable $x_{jm}$, also of size $p_{j}$, that represents the option of processing job $j$ on machine $m\in\mathcal{M}$. The presence of the interval $x_{jm}$ is optional since it is allowed to take the value $\perp$, which indicates its absence in the solution. These optional interval variables allow the model to decide which machine should process the jobs. Equation (1) defines a function ${\mathfrak{f}}_{m}$, which is the unique state of the machine $m$ at every point in time. Equation (1) defines a cumulative function ${\mathscr{l}}_{mf}$, which represents the load on machine $m$ of jobs of the family $f\in\mathcal{F}$. It uses the pulse function, which receives two inputs: An interval $x_{jm}$ and an integer value $s_{j}$. This pulse function tells the cumulative function to increment its value in $s_{j}$ units during the interval $x_{jm}$, if present. Hence, the function ${\mathscr{l}}_{mf}$ accumulates the sizes of overlapping jobs of family $f\in\mathcal{F}$ on machine $m$ at every point in time.
@@
$x_{j}$$\in\{[a,a+p_{j}):a\in\mathbb{Z},r_{j}\leq a\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J};$ $(1)$
@@ VI-B The Synchronized Model
$x_{jm}^{b}$$\in\{[a,a+p_{j}:r_{j}\leq a\in\mathbb{Z})\}\cup\{\perp\},$
$\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M},b\in\mathcal{B}_{f_{j}};$ $(22)$
@@
Model 3 presents the Synchronized (S) model. It is based on the CP model by Cakici et al. [10] for AM processes, which deals with a single compatible family; this paper extends their formulation to handle multiple incompatible families. Equation (3) defines an optional interval variable $x_{jm}^{b}$, of size $p_{j}$, that represents the option of job $j$ being batched in batch $b\in\mathcal{B}_{f_{j}}$ and processed on machine $m\in\mathcal{M}$. Equation (3) defines an optional interval variable $y_{bm}$ of size $p^{f^{b}}$ that represents the option of batch $b\in\mathcal{B}$ being processed on machine $m$. Equation (3) defines an interval sequence variable $\varphi_{m}$ that represents the sequence of batches scheduled on machine $m$. Equation (3) defines a cumulative function ${\mathscr{l}}_{m}^{b}$ that represents the load on machine $m$ due to the jobs in batch $b$.
@@
Objective function (3) minimizes the TWCT. Constraints (3) assign each job to exactly one batch on exactly one machine. Constraints (3) allow each batch to be processed in at most one machine. Constraints (3) use the synchronize global constraint, which takes two inputs: an interval variable $v=x_{jm}^{b}$ and a set of interval variables $B=\set{y_{bm}}$. If $v$ is present, this constraint attempts to synchronize it with any intervals in $B$ that are also present [38], i.e., $\texttt{presenceOf}(v)=1\implies\texttt{startOf}(v)=\texttt{startOf}(b)\wedge\texttt{endOf}(v)=\texttt{endOf}(b),\forall\leavevmode\nobreak\ b\in B$ s.t. $\texttt{presenceOf}(b)=1$. In practice, when the interval variable representing the job being assigned to a batch on a machine (i.e., $x_{jm}^{b}$) is part of the solution, the synchronize constraint attempts to align its start and end times with those of the batch on the machine (i.e., $y_{bm}$), if it is also present. However, the synchronize constraint only ensures synchronization if both interval variables are present; it does not inherently enforce the presence of $y_{bm}$ when $x_{jm}^{b}$ is scheduled. This subtle issue, where the presence of $x_{jm}^{b}$ should logically imply the presence of $y_{bm}$ is addressed in the original model by Cakici et al. [10] by means of an additional constraints that defines the correct duration of the batches depending on the jobs assigned to it, if any. This constraint is necessary in their AM application as the duration of the batch depends on the size of the parts 3D-printed in it. However, this constraint is not required in the context of the diffusion area studied in this paper since the duration of the batches is already predefined. Nonetheless, the presence of the correct batches is still needed to be enforced. To resolve this, constraints (3) introduce an explicit implication constraint which ensures that, whenever $x_{jm}^{b}$ is included in the solution, the corresponding batch interval $y_{bm}$ is also present. Constraints (3) use the noOverlap global constraint to ensure that the present interval variables in the machine sequence do not overlap, guaranteeing at most one batch being processed at a time on each machine. Finally, constraints (3) guarantee that the load of the machines when processing each job in its batch is within the size limits of the batch’s family.
@@ VI-C The Redundant Synchronized Model
$x_{j}$$\in\{[a,a+p_{j}:r_{j}\leq a\in\mathbb{Z})\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J};$ $(33)$
$x_{jm}$$\in\{[a,a+p_{j}:a\in\mathbb{Z})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(34)$
$x_{jm}^{b}$$\in\{[a,a+p_{j}:a\in\mathbb{Z})\}\cup\{\perp\},$
$\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M},b\in\mathcal{B}_{f_{j}};$ $(35)$
$y_{b}$$\in\{[a,a+p^{f^{b}}:a\in\mathbb{Z})\}\cup\{\perp\},\nobreak\leavevmode$ $\forall\leavevmode\nobreak\ b\in\mathcal{B};$ $(36)$
$y_{bm}$$\in\{[a,a+p^{f^{b}}:a\in\mathbb{Z})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(37)$
@@
$V_{jm}^{b}$ =$=\{y_{b},y_{bm},x_{jm}\},\quad\quad\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M},b\in\mathcal{B}_{f_{j}};$ $(39)$
@@ VIII Concluding remarks
This paper addressed the incompatible case of parallel batch scheduling, where compatible jobs belong to the same family, and jobs of different families cannot be processed together in the same batch. In the context of the diffusion area of the SM process, batches cannot be interrupted mid-process and jobs must be synchronized within their batch. The NS model by Ham et al. [18] for this problem does not synchronize the jobs within their batch, allowing interruptions mid-process. To remedy this limitation, the paper presented three new CP models, including two adaptions of existing models and a novel RS model. The AS model of Nascimento et al. [8] was extended to handle multiple incompatible job families and satisfy the batch size requirements. The S model of Cakici et al. [10] was also extended to the incompatible case. Finally, the paper proposed a new RS model that exclusively uses global constraints and adds redundancy to the problem structure. The computational experiments demonstrate that the proposed RS model dominates all the other models: it finds solutions to all the instances, proves optimality in the largest amount of instances, produces solutions that are significantly better than other models, especially in the largest instances and achieves the best gaps in the least amount of time. Current research includes using CP for serial batching, and future research includes extending the efficient arc-flow formulations for compatible p-batch [16, 29, 28, 30] to the incompatible case and embedding the proposed formulations in larger scheduling contexts [40, 41]